# RSTeller: Scaling Up Visual Language Modeling in Remote Sensing with Rich Linguistic Semantics from Openly Available Data and Large Language Models

### v1

@@ 2.2.1 Descriptive Tasks
Prompt assembly: Once the attributes and tags have been interpreted, the final step in Task 1 is to assemble all the interpretations into a coherent prompt for the LLM. This involves organizing the interpreted attributes and tag descriptions into a structured format that the LLM can effectively process. In this work, we employ a template-based approach for prompt assembly, where we define a task-specific template to structure the prompt and fill in the placeholders with the interpreted attributes and tags. A brief showcase of the prompt template for Task 1 is presented in Fig. 3. The overall structure of the prompt is divided into three parts. First, general instructions are provided to the LLM under the headings ###Instruction### and ###Captioning Objective###. The former offers general task instructions and some basic knowledge about the attributes used, while the latter highlights the attributes crucial for the task and details that require particular attention. Specifically, we instruct the LLM to not only describe the selected element but also infer its possible surroundings and context based on the provided information. The inferred contexts by the LLM should be highlighted by some words denoting uncertainty, such as ”possible” or ”may”. This encourages the LLM to generate richer and more informative captions, expanding the scope of the task beyond the selected element to the entire image patch. Next, a one-shot example is included to illustrate the format and style of the desired output, helping the LLM understand the task requirements more clearly. The example follows the same format as the real task, consisting of a ”Raw” and ”Caption” pair. The ”Raw” part includes attributes and tags interpreted from an example image patch, while the ”Caption” part is the gold-standard caption from a human annotator and revised by GPT-4. Lastly, task-specific inputs are provided. The ”Raw” part is a predefined context with multiple placeholders, where the interpreted attributes and tags are directly filled, except for the ”Partition” attribute. If the ”Partition” attribute is True, the placeholder is replaced with the sentence ”Some parts of the geometry extended out of this ROI,” otherwise, it is left blank. The ”Caption” part remains blank, awaiting the LLM’s continual writing. This assembly step ensures that the LLM’s output is both precise and aligned with the intended use of the remote sensing data, facilitating effective and insightful captioning.
@@
Task 2 (Non-area description): Similar to Task 1, this task involves four steps. Based on our inspection of the collected OSM data, the ”non-area” with the longest length is sometimes a simple line from a large geometry, which is not well annotated by OSM tags. Therefore, unlike Task 1, which selects the OSM element with the largest size, Task 2 first identifies three OSM elements with the longest length and then select the one with the most tags as the most distinctive element to be described.
In the attribute interpretation step, we extract different attributes compared to Task 1 due to the geometrical differences between the ”area” and ”non-area” elements. The attributes for Task 2 are shown in Table 3. To better describe the position of the non-area element, we use endpoint locations instead of the coarse location derived from the element’s center. The nine-grid system, the same as used in Task 1, assigns labels to the endpoints. The sinuosity attribute describes the shape of the element, with categories including ”straight,” ”curved,” ”twisted,” ”closed,” and ”broken,” determined by thresholding the ratio between the element’s length and the distance between its endpoints. The normalized length attribute represents the relative length of the element compared to the square root of the image patch area. The length attribute is the actual length of the element in meters. The orientation attribute, determined by calculating the angle between the endpoints and the horizontal axis, describes the direction of the element with categories like ”W_E,” ”S_N,” ”SW_NE,” and ”NW_SE.” The simplified geometry attribute represents the simplified and normalized geometry of the element, following the same scheme as in Task 1. Finally, the is-cropped attribute indicates whether the element is cropped by the image patch.
@@
Endpoint Locations ”(bottom-center, right-center)”, ”(left-bottom, right-bottom)”, etc.
Sinuosity ”straight”, ”curved”, ”twisted”, ”closed”, ”broken”
Normalized Length ”0.351”, ”1.260”, etc.
Length ”94”, ”339”, etc.
Orientation ”W_E”, ”S_N”, ”SW_NE”, ”NW_SE”
Simplified Geometry ”[(0.211, 0.000), (0.320, 0.794), …]”, etc.
@@ 2.2.2 Caption Augmentation Task
The caption augmentation task aims to enhance the diversity and quality of the initial captions generated for the image patches, creating multiple revisions from the raw caption with different tones. This task, denoted as Task 3 in this work and illustrated in Fig. 2, is relatively simple compared to previous tasks involving OSM element selection and interpretation, as it is purely a linguistic task and does not require specialized knowledge. Consequently, it comprises only two steps: revision example sampling and prompt assembly.
Unlike previous tasks that utilized a fixed example for each task, this task necessitates a larger number of examples to produce diverse and informative captions. To this end, we constructed a meta-example set containing multiple raw captions and their corresponding revisions. In this work, we separately selected five raw captions from Task 1 and Task 2 and generated five revisions for each raw caption, altering the tone, inflection, word choice, and length of the original caption. These revisions were written by human annotators and polished by GPT-4. In Task 3, examples are dynamically sampled from the meta-example set, selecting all five raw captions of the corresponding task and randomly choosing one of the five revisions for each raw caption. The five raw and revision pairs are then shuffled and used as the five-shot examples for the prompt assembly step. This approach ensures that the revisions generated by the LLM are diverse and informative.
In the prompt assembly step, a simpler prompt template is utilized, as shown in Fig. 4. The overall structure of the prompt template is similar to that of Tasks 1 and 2, divided into three parts: general instructions, few-shot examples, and task inputs. As the revision task is straightforward and does not require detailed caption objectives or specialized knowledge about the captioned element’s attributes, the general instructions are kept concise, informing the LLM to revise an existing caption, specifying the coordinate system used, and indicating the provided examples. The five-shot examples in the prompt template are defined by several placeholders, filled by sampled examples from the previous step. The task inputs include a placeholder for the caption to be revised, leaving the revised result empty for the LLM to fill in.
@@ 3 RSTeller Dataset
Utilizing the proposed dataset generation workflow, we have developed the RSTeller dataset. The data acquisition and processing were conducted using a local server equipped with 128 cores, 512 GB of memory, and four NVIDIA A800 GPUs. We deployed one Mixtral-7B instance on each GPU to generate the captions in parallel. The data acquisition process took approximately four months, during which we collected roughly 15 million image patches. Of these, approximately 7 million had corresponding OSM data fetched from the Kumi endpoint. The captioning process ran concurrently with data acquisition and processing, achieving a maximum throughput of 50,000 captions per day. This process spanned over two months and generated approximately 3 million captions. Ultimately, 2.5 million high-quality image-text pairs were selected and compiled into the RSTeller dataset. The following sections provide detailed insights into the dataset.
@@ 3.1 Dataset Overview
RSTeller dataset comprises approximately 1.2 million remote-sensing image patches, each accompanied by multiple captions, ranging from 2 to 5, resulting in roughly 2.5 million image-text pairs. The images are sourced from the National Agriculture Imagery Program (NAIP) on the GEE platform, with a ground sample distance (GSD) of 0.6 meters. Geographically, the dataset focuses on the United States, with images acquired via aircraft. Detailed attributes of the dataset are presented in Table 4.
@@
Number of Images 1,197,190
Number of Image-Text Pairs 2,539,256
@@
The distribution of captions within the dataset is illustrated in Fig. 5. As shown in Fig. 5(a), Task 2 captions slightly outnumber Task 1 captions, while Task 3 captions constitute more than half of the entire dataset. The prevalence of Task 3 captions is due to the multiple revisions of the raw captions. Fig. 5(b) indicates that while most image patches (88.44%) have two caption variations (one raw caption and one revision), some images have up to five variations, enhancing the dataset’s diversity in linguistic variations.
Fig. 5: Distribution of captions in the RSTeller dataset. (a) Proportions of captions categorized by tasks. (b) Relationship between the number of images and caption variations.
@@
To facilitate qualitative analysis, we have provided a sample inspection of the dataset in Fig. 7. This figure presents twenty image patches along with their corresponding OSM information and captions, sampled from various locations across the United States. The map shows red dots indicating the sample locations, with a digit labeling each patch ID. A topographic base map from OpenTopoMap [^8] https://opentopomap.org is included to provide a visual reference for the geographical locations of the samples. The figure then displays the image patches, selected OSM elements, their tags, and the captions under specific captioning tasks. Selected OSM elements are highlighted in red: area elements are highlighted with masks and enclosed by their respective boundaries, while non-area elements are highlighted with simple lines.
@@
Fig. 7: Inspection of the RSTeller dataset.
Fig. 8: Inspection of the RSTeller dataset (continued).
@@ 3.2 Caption Analysis
First, we examine the distribution of caption lengths, determined by counting the number of tokens using the Natural Language Toolkit (NLTK) word tokenizer. The caption lengths vary from 5 to 192 tokens, with a median length of 48 tokens and an average length of 54.2 tokens and its distribution is depicted in Fig. 9(a). This variation underscores the diversity within the captions, which could enhance the robustness of VLM training. However, given that the CLIP model has a maximum input length of 77 tokens, captions exceeding this limit may compromise performance, highlighting the need for meticulous preprocessing.
Moreover, we assess the metadata richness by analyzing the number of tags associated with each OSM element used for captioning. As Fig. 9(b) illustrates, while most images feature fewer than 10 tags, some are annotated with over 30, with a maximum of 63, indicating a skewed distribution with a long tail.
@@
Subsequently, we scrutinize the most frequent non-stop words in the captions to assess content richness. The top 50 most frequent words are shown in Fig. 10. Common but meaningless stop words such as ”the”, ”is”, and ”and” are filtered out. The remaining words are sorted by their frequency in descending order. Words suggesting inferential reasoning like ”likely” and ”possibly” are prevalent, reflecting the LLM’s inference from the OSM tags and its knowledge. For example, in image patch 8 of Fig. 7, the OSM tag only contains the tag ”landuse: farmland,” but the caption generated by the LLM extends the semantics by adding ”It is likely used for growing crops,” connecting shallow tag information to deeper meanings. Additionally, words like ”meters,” ”within,” and ”orientation” appear frequently, demonstrating the richness of shape and geography information in the dataset due to the interpretation of attributes in the captioning tasks. For instance, the caption for image patch 3 in Fig. 7 precisely describes the road’s orientation (south to north) and length (270 meters).
@@
Additionally, we delve into the topical diversity of the captions by investigating the frequency of the top 20 keys and their corresponding values within the OSM tags used for captioning. As shown in Fig. 11(a), the top 20 most frequent keys include feature-related keys such as highway, natural, landuse, waterway, and building, while others are more general like name and TIGER (Topologically Integrated Geographic Encoding and Referencing), which are usually unique identifiers. To further explore the attribute distribution of some features, we survey the top 20 most frequent values from keys highly related to image features. Fig. 11(b-f) display these for highway, natural, landuse, waterway, and surface. All values exhibit long-tailed distributions, with significant differences in frequency. Notably, even the most frequent key ”highway” has a frequency of no more than 500k, which is less than half of the total number of OSM elements used for captioning the entire dataset, indicating a diverse and rich set of attributes used for captioning.
@@
In conclusion, while the captions in the RSTeller dataset are rich and diversified, focusing on varied remote sensing features, they also contain ambiguities and inferred content that may introduce noise. These attributes are pivotal for training sophisticated VLMs in the remote sensing domain, but they also necessitate careful handling to mitigate potential biases and ensure the utility of the trained models.

### v4

@@ 2.2.1 Descriptive Tasks
Prompt assembly: Once the attributes and tags have been interpreted, the final step in Task 1 is to assemble all the interpretations into a coherent prompt for the LLM. This involves organizing the interpreted attributes and tag descriptions into a structured format that the LLM can effectively process. In this work, we employ a template-based approach for prompt assembly, where we define a task-specific template to structure the prompt and fill in the placeholders with the interpreted attributes and tags. A brief showcase of the prompt template for Task 1 is presented in Fig. 4. The overall structure of the prompt is divided into three parts. First, general instructions are provided to the LLM, offering a high-level overview of the task followed by detailed guidance on completing it. Specifically, the instructions emphasize producing a fluent and natural description of the selected element. The LLM is instructed to describe the element based on its visual attributes, such as position, shape, approximate size, and notable features, while incorporating reasonable inferences about the surrounding context using cautious language (e.g., “possible,” “likely”). This approach expands the task scope beyond the selected element to include the broader scene. Next, a five-shot example is included to illustrate the format and style of the desired output, helping the LLM understand the task requirements more clearly. The example follows the same format as the real task, consisting of a “Raw” and “Caption” pair. The “Raw” part includes attributes and tags interpreted from an example image patch, while the “Caption” part is the gold-standard caption from a human annotator and revised by GPT-4. The human annotator ensures that the caption captures domain-specific details accurately, but minor issues such as spelling, phrasing, or grammar errors can still occur. GPT-4 is then used to refine the caption, improving its fluency, consistency, and overall quality while maintaining its original meaning. Finally, the revised caption is double-checked by the annotator to confirm its correctness and domain alignment. This combined approach leverages both human expertise and GPT-4’s linguistic capabilities to produce high-quality captions. Lastly, task-specific inputs are provided. The “Raw” part is a predefined context with multiple placeholders, where the interpreted attributes and tags are directly filled, except for the “is-cropped” attribute. If the “is-cropped” attribute is True, the placeholder is replaced with the sentence “Some parts of the geometry extend beyond this ROI,” otherwise, it is left blank. The “Caption” part remains blank, awaiting the LLM’s continual writing. This assembly step ensures that the LLM’s output is both precise and aligned with the intended use of the remote sensing data, facilitating effective and insightful captioning. The complete prompts for this task, as well as others used in this work, are publicly available in our GitHub repository for transparency and reproducibility.
@@
Task 2 (Non-area description): Similar to Task 1, this task involves four steps. The element selection for Task 2 follows a similar structure to Task 1 but with different criteria. We first filter out elements whose length is smaller than 30% of the square root of the image patch area. From the remaining elements, we randomly select one from the top three longest.
In the attribute interpretation step, we extract different attributes compared to Task 1 due to the geometrical differences between the “area” and “non-area” elements. The attributes for Task 2 are shown in Table 3. Endpoint locations replace the coarse location for non-area elements, capturing positional information specific to linear features. This ensures accurate spatial descriptions for elongated structures like roads or rivers. The nine-grid system, the same as used in Task 1, assigns labels to the endpoints. The sinuosity attribute captures the geometric complexity of linear elements, offering insights into their curvature and form, which are crucial for distinguishing between infrastructure types like roads or waterways. Its categories including “straight,” “curved,” “twisted,” “closed,” and “broken,” determined by thresholding the ratio between the element’s length and the distance between its endpoints. The normalized length attribute represents the relative length of the element compared to the square root of the image patch area, enabling models to assess the spatial extent of linear features relative to the image context. The length attribute is the actual length of the element in meters, offering precise quantitative information that enhances spatial and contextual reasoning. The orientation attribute, determined by calculating the angle between the endpoints and the horizontal axis, describes the direction of the element with categories like “west-east,” “south-north,” “southwest-northeast,” and “northwest-southeast”, aiding the model in understanding spatial relationships and movement patterns within the image patch. For non-area elements exhibiting extreme curvature with a sinuosity value exceeding 1.5 (the defined threshold for distinguishing curved elements), the orientation is labeled as “too curved or twisted to determine accurately.” This ensures that the attribute remains meaningful and robust, even in cases of complex geometries. The simplified geometry attribute represents the simplified and normalized geometry of the element, following the same scheme as in Task 1. Finally, the is-cropped attribute indicates whether the element is cropped by the image patch. Given the differences in the nature of “area” and “non-area” elements, as well as the distinct attributes and their interpretation processes, we use separate icons to emphasize these variations in Fig. 2. This differentiation enhances clarity and facilitates a better understanding of the respective workflows.
In some cases, a non-area element may also consist of separated lines, similar to the behavior in area elements. In such instances, we interpret the endpoint locations, sinuosity, and orientation based solely on the longest line within the element, while other attributes are derived from the entire element as a whole.
@@
Endpoint Locations “(bottom-center, right-center)”, “(left-bottom, right-bottom)”, etc.
Sinuosity “straight”, “curved”, “twisted”, “closed”, “broken”
Normalized Length “0.351”, “1.260”, etc.
Length “94”, “339”, etc.
Orientation ‘west-east”, “south-north”, etc.
Simplified Geometry “[(0.211, 0.000), (0.320, 0.794), …]”, etc.
@@ 2.2.2 Caption Augmentation Task
The caption augmentation task aims to enhance the diversity and quality of the initial captions generated for the image patches, creating a revision from the raw caption with different tones. This task, denoted as Task 3 in this work and illustrated in Fig. 2, is relatively simple compared to previous tasks involving OSM element selection and interpretation, as it is purely a linguistic task and does not require specialized knowledge. Consequently, it comprises only two steps: revision example sampling and prompt assembly.
Unlike previous tasks that utilized a fixed example for each task, this task necessitates a larger number of examples to produce diverse and informative captions. To this end, we constructed a meta-example set containing multiple raw captions and their corresponding revisions. In this work, we separately selected five raw captions from Task 1 and Task 2 and generated five revisions for each raw caption, altering the tone, inflection, word choice, and length of the original caption. These revisions were written by human annotators and polished by GPT-4 following the same approach as in Task 1 and Task 2. In Task 3, examples are dynamically sampled from the meta-example set, selecting all five raw captions of the corresponding task and randomly choosing one of the five revisions for each raw caption. The five raw and revision pairs are then shuffled and used as the five-shot examples for the prompt assembly step. This approach ensures that the revisions generated by the LLM are diverse and informative.
In the prompt assembly step, a prompt template is utilized, as shown in Fig. 5. The overall structure of the prompt template is similar to that of Tasks 1 and 2, divided into three parts: general instructions, few-shot examples, and task inputs. The general instructions are crafted to guide the LLM in revising existing captions, emphasizing variations in tone, phrasing, and length while maintaining the core meaning. These instructions also provide clarity on the coordinate system used and highlight the role of examples in facilitating the task. The five-shot examples in the prompt template are defined by several placeholders, filled by sampled examples from the previous step. The task inputs include a placeholder for the caption to be revised, leaving the revised result empty for the LLM to fill in.
@@ 3 RSTeller Dataset
Utilizing the proposed dataset generation workflow, we have developed the RSTeller dataset. The data acquisition and processing were conducted using a local server equipped with 128 cores, 512 GB of memory, and four NVIDIA A800 GPUs. To generate captions efficiently, we utilized two instances of Mixtral-Nemo-Instruct and two instances of Mixtral-Small-Instruct, each assigned to a single GPU. The captioning tasks were dynamically distributed based on GPU occupancy to maximize parallel processing efficiency. To accelerate the captioning process, the vLLM framework (Kwon et al., 2023) was utilized to host the large language models. The data acquisition process took approximately four months, during which we collected roughly 15 million image patches. Of these, approximately 7 million had corresponding OSM data fetched from the Kumi endpoint. The captioning process ran concurrently with data acquisition and processing, achieving an average throughput of about 200 thousand captions per day. This process spanned over two weeks and generated approximately 3 million captions. Ultimately, 2.6 million well-crafted image-text pairs were selected and compiled into the RSTeller dataset. The following sections provide detailed insights into the dataset.
@@ 3.1 Dataset Overview
RSTeller dataset comprises approximately 1.3 million remote-sensing image patches, each accompanied by two captions, resulting in roughly 2.6 million image-text pairs. The images are sourced from the National Agriculture Imagery Program (NAIP) on the GEE platform, with a ground sample distance (GSD) of 0.6 meters. Geographically, the dataset focuses on the United States, with images acquired via aircraft. Detailed attributes of the dataset are presented in Table Table 4.
@@
Number of Images 1,309,926
Number of Image-Text Pairs 2,619,852
@@
Fig. 6 illustrates the distribution of captions across tasks in the RSTeller dataset. Task 1 captions (area elements) are slightly fewer than Task 2 captions (non-area elements). Task 3 captions, comprising revised versions of the raw captions, equal the combined total of Task 1 and Task 2 captions. Shaded regions indicate annotations generated by Mixtral-Small-Instruct, while solid regions correspond to annotations from Mixtral-Nemo-Instruct. The lower output of Mixtral-Small-Instruct is attributed to its larger model size and reduced overall throughput.
Fig. 6: Distribution of captions across tasks in the RSTeller dataset. Solid colors indicate annotations generated by Mixtral-Nemo-Instruct, while shaded areas represent annotations from Mixtral-Small-Instruct.
@@
To facilitate qualitative analysis, we have provided a sample inspection of the dataset in Fig. 3. This figure presents four image patches along with their corresponding OSM information and captions, sampled from various locations across the United States. Topographic base maps from OpenTopoMap [^8] https://opentopomap.org are included to provide a visual reference for the geographical locations of the samples. The figure then displays the image patches, selected OSM elements, their tags, and the captions under specific captioning tasks. Selected OSM elements are highlighted in red: area elements are highlighted with masks and enclosed by their respective boundaries, while non-area elements are highlighted with simple lines. OSM tags not incorporated into the captioning process are crossed out.
@@ 3.2 Caption Analysis
First, we examine the distribution of caption lengths, determined by counting the number of tokens using the Natural Language Toolkit (NLTK) word tokenizer. The caption lengths vary from 2 to 155 tokens, with a median length of 42 tokens and an average length of 40.9 tokens and its distribution is depicted in Fig. 8(a). This variation underscores the diversity within the captions, which could enhance the robustness of VLM training. However, given that the CLIP model has a maximum input length of 77 tokens, captions exceeding this limit may compromise performance, highlighting the need for meticulous preprocessing.
We further evaluate metadata richness by analyzing the number of tags associated with each OSM element used for captioning. As shown in Fig. 8(b), the distribution displays a bimodal pattern, with a prominent peak in the 0-10 tags range, a secondary peak around 45-50 tags, and a noticeable dip between 25 and 35 tags. The secondary peak primarily arises from OSM elements related to protected areas, which often include tag groups starting with prefixes such as ”nconemap” or ”NCOS.” These tag groups, encompassing attributes like land ownership, conservation management, and access restrictions, typically contain extensive metadata relevant to image captioning. As such, they are retained during preprocessing to preserve valuable information.
@@
[…]
@@
Additionally, we delve into the topical diversity of the captions by investigating the frequency of the top 20 keys and their corresponding values within the OSM tags used for captioning all the image patches. As shown in Fig. 10(a), the top 20 most frequent keys include feature-related keys such as “highway”, “natural” and “landuse”, while others are more general like “name” and “tiger:county” which are usually unique identifiers. To further explore the attribute distribution of some features, we survey the top 20 most frequent values from keys highly related to image features. Fig. 10(b-f) display these for “highway,” “natural,” “landuse,” “waterway,” and “surface”. All values exhibit long-tailed distributions, with significant differences in frequency. The richness of the tag variety ensures the diversity in the topics of all the captions.
@@
Finally, we assess the overall semantic richness of the captions by calculating the MTLD score (McCarthy and Jarvis, 2010), a widely recognized metric for evaluating linguistic diversity (Ding et al., 2023; Herbold et al., 2023; Bain et al., 2021), and compare the results with previous works. The MTLD score is computed by treating all captions in the dataset as a single document. To minimize bias introduced by multiple captions for the same image, the captions are concatenated in a random order before calculating the MTLD score for the resulting string. The MTLD score is defined as the mean length of sequential word segments in the text that maintain a minimum threshold Type-Token Ratio (TTR), where TTR represents the ratio of unique words (types) to the total number of words (tokens) in the text, denoted as $\text{TTR}_{\text{thr}}$. This process is mathematically expressed in Eq. 1.
MTLD=2×Total $\text{MTLD}=\frac{2\times\text{Total Words}}{\text{Factor}_{\text{fwd}}+\text{Factor}_{\text{bwd}}},$ (1)
where Total Words denotes the total token count of the concatenated text, and Factor represents the number of segments where each segment maintains a TTR no lower than the predefined threshold, computed both forward ($\text{Factor}_{\text{fwd}}$) and backward ($\text{Factor}_{\text{bwd}}$).
Segmentation begins with the first word, and the cumulative TTR is computed as new words are added. The process halts when the TTR drops below the threshold, forming a segment. The process then restarts from the next word, iterating until the entire text is processed. If a segment is incomplete at the end of the text, a partial factor is computed as $\frac{1-\text{TTR}}{1-\text{TTR}_{\text{thr}}}$. For this evaluation, a default TTR threshold of 0.72 is used. The MTLD score quantifies lexical diversity by assessing how consistently unique words are introduced throughout the dataset. Higher MTLD scores indicate greater lexical diversity, reflecting richer vocabulary usage and less repetitive patterns across captions. The results are illustrated in Fig. 1.
As shown in Fig. 1, RSTeller achieves an MTLD score exceeding 100, indicating the highest semantic richness among the datasets. In comparison, earlier datasets such as Sydney-Captions, RSICD, and NWPU-Captions exhibit lower MTLD scores and much smaller dataset scales, underscoring their limitations in both semantic richness and scale. More recent datasets, such as RS5M and SkyScript, achieve moderate improvements in scale but fail to match the semantic diversity and vocabulary richness of RSTeller. Notably, SkyScript has the lowest MTLD score due to its rigid heuristic approach for generating captions. All captions in SkyScript follow a repetitive pattern, such as “a satellite image of XX of XX,” which significantly limits the linguistic diversity and is a major factor contributing to its low score.
In conclusion, the captions in the RSTeller dataset are not only rich and diverse, capturing a wide range of remote sensing features, but also substantial in scale. While these captions include ambiguities and inferred content that may introduce noise, the richness, diversity, and scale of the data are critical for training advanced VLMs in the remote sensing domain. This is particularly true during the pre-training stage, where large-scale data is essential for acquiring foundational vision-language knowledge, and minor levels of noise can be tolerated. However, careful management of these attributes is necessary to minimize potential biases and ensure the reliability and effectiveness of the trained models.